OpenStudy (anonymous):

I need help! I tried but got it wrong. Write the equation of a line that passes through the point (3, 1) and is parallel to the line y = x + 2.

OpenStudy (anonymous):

Y=5/3+2

OpenStudy (lgbasallote):

the slope of the equation y = x + 2 is 1 because the coefficient of x is 1 therefore the slope of the parallel line is also 1 so now just use the point-slope formula \[y - y_1 = m(x-x_1)\] in this case x1 is 3 because that's the x coordinate; y1 is 1 because that's the y coordinate; m is 1 because like i said earlier the slope is 1 \[\implies y - 1 = 1(x-3)\] distribute 1 into x-3 \[\implies y - 1 = x-3\] add 1 to both sides \[\implies y - 1 + 1 = x - 3+1\] combine like terms \[\implies y = x-2\] did you get that?

OpenStudy (lgbasallote):

feel free to ask anything about what i did

OpenStudy (anonymous):

Its actually y=5/3x+2 not y=x+2. I'm sorry